OpenStudy (gavin39):

Ramona is playing a board game and rolls two number cubes. Let A = {the sum of the number cubes is odd} and let B = {the sum of the number cubes is divisible by 5}. List the outcomes in A ∩ B. {1,3,5,7,9,11} {1,5} {3,5} {5}

OpenStudy (gavin39):

@minoz

OpenStudy (anonymous):

A=1,3,5,7,9,11 as it is odd 2nd is 5,10 as only 5,10 is ddivisible by 5 intersection is common so only 5 appears

OpenStudy (gavin39):

.What are the amplitude and midline? amplitude: 1; midline: y = 1 amplitude: 0; midline: y = 0 amplitude: 2; midline: y = 1 amplitude: 2; midline: y = 0

OpenStudy (anonymous):

midline=(2+0)/2 amplitude=2-mid line

OpenStudy (gavin39):

So this one is C, right?

OpenStudy (anonymous):

no A amp=2-1